\begin{thebibliography}{}

\bibitem[Allam and Haggag, 2012]{allam2012question}
Allam, A. M.~N. and Haggag, M.~H. (2012).
\newblock {The Question Answering Systems: A Survey}.
\newblock {\em International Journal of Research and Reviews in Information
  Sciences (IJRRIS)}, 2(3).

\bibitem[Ammar et~al., 2020]{ammar2020methodology}
Ammar, A., Mehryar, S., and Celebi, R. (2020).
\newblock A methodology for hierarchical classification of semantic answer
  types of questions.
\newblock In {\em Proceedings of the SeMantic AnsweR Type prediction task
  (SMART 2020)}, pages 41--48.

\bibitem[Azmy et~al., 2018]{azmy2018farewell}
Azmy, M., Shi, P., Lin, J., and Ilyas, I. (2018).
\newblock {Farewell freebase: Migrating the Simple Questions dataset to
  DBpedia}.
\newblock In {\em Proceedings of the 27th International Conference on
  Computational Linguistics}, pages 2093--2103.

\bibitem[Bill and Jim{\'e}nez-Ruiz, 2020]{bill2020question}
Bill, E. and Jim{\'e}nez-Ruiz, E. (2020).
\newblock Question embeddings for semantic answer type prediction.
\newblock In {\em Proceedings of the SeMantic AnsweR Type prediction task
  (SMART 2020)}, pages 71–--80.

\bibitem[Diefenbach et~al., 2017]{wikidata-benchmark}
Diefenbach, D., Tanon, T.~P., Singh, K.~D., and Maret, P. (2017).
\newblock Question answering benchmarks for wikidata.
\newblock In {\em Proceedings of the {ISWC} 2017 Posters {\&} Demonstrations
  and Industry Tracks co-located with 16th International Semantic Web
  Conference {(ISWC} 2017), Vienna, Austria, October 23rd - to - 25th, 2017.}

\bibitem[Dubey et~al., 2019]{dubey2019lc}
Dubey, M., Banerjee, D., Abdelkawi, A., and Lehmann, J. (2019).
\newblock {LC-QuAD 2.0: A Large Dataset for Complex Question Answering over
  Wikidata and DBpedia}.
\newblock In {\em International Semantic Web Conference}, pages 69--78.
  Springer.

\bibitem[Dubey et~al., 2018]{earl}
Dubey, M., Banerjee, D., Chaudhuri, D., and Lehmann, J. (2018).
\newblock Earl: joint entity and relation linking for question answering over
  knowledge graphs.
\newblock In {\em ISWC2018}, pages 108--126.

\bibitem[Harabagiu et~al., 2000]{harabagiu2000falcon}
Harabagiu, S., Moldovan, D., Pasca, M., Mihalcea, R., Surdeanu, M., Bunescu,
  R., Girju, R., Rus, V., and Morarescu, P. (2000).
\newblock Falcon: Boosting knowledge for answer engines.
\newblock In {\em TREC}, volume~9, pages 479--488.

\bibitem[Kertkeidkachorn et~al., 2020]{kertkeidkachorn2020hierarchical}
Kertkeidkachorn, N., Nararatwong, R., Nguyen, P., Yamada, I., Takeda, H., and
  Ichise, R. (2020).
\newblock Hierarchical contextualized representation models for answer type
  prediction.
\newblock In {\em Proceedings of the SeMantic AnsweR Type prediction task
  (SMART 2020)}, pages 49--56.

\bibitem[Li and Roth, 2006]{li2006learning}
Li, X. and Roth, D. (2006).
\newblock {Learning Question Classifiers: the Role of Semantic Information}.
\newblock {\em Natural Language Engineering}, 12(3):229--249.

\bibitem[Lin et~al., 2020]{lin2020kbpearl}
Lin, X., Li, H., Xin, H., Li, Z., and Chen, L. (2020).
\newblock Kbpearl: a knowledge base population system supported by joint entity
  and relation linking.
\newblock {\em Proceedings of the VLDB Endowment}, 13(7):1035--1049.

\bibitem[Mihindukulasooriya et~al., 2020a]{DBLP:conf/semweb/2020smart}
Mihindukulasooriya, N., Dubey, M., Gliozzo, A., Lehmann, J., Ngomo, A.~N., and
  Usbeck, R., editors (2020a).
\newblock {\em Proceedings of the SeMantic AnsweR Type prediction task
  {(SMART)} at {ISWC} 2020 Semantic Web Challenge co-located with the 19th
  International Semantic Web Conference {(ISWC} 2020), Virtual Conference,
  November 5th, 2020}, volume 2774 of {\em {CEUR} Workshop Proceedings}.
  CEUR-WS.org.

\bibitem[Mihindukulasooriya et~al., 2020b]{mihindukulasooriya2020leveraging}
Mihindukulasooriya, N., Rossiello, G., Kapanipathi, P., Abdelaziz, I.,
  Ravishankar, S., Yu, M., Gliozzo, A., Roukos, S., and Gray, A. (2020b).
\newblock Leveraging semantic parsing for relation linking over knowledge
  bases.
\newblock In {\em International Semantic Web Conference}, pages 402--419.
  Springer.

\bibitem[Naseem et~al., 2021]{naseem2021semantics}
Naseem, T., Ravishankar, S., Mihindukulasooriya, N., Abdelaziz, I., Lee, Y.-S.,
  Kapanipathi, P., Roukos, S., Gliozzo, A., and Gray, A. (2021).
\newblock A semantics-aware transformer model of relation linking for knowledge
  base question answering.
\newblock In {\em Proceedings of the 59th Annual Meeting of the Association for
  Computational Linguistics and the 11th International Joint Conference on
  Natural Language Processing (Volume 2: Short Papers)}, pages 256--262.

\bibitem[Ngomo, 2018]{ngomo20189th}
Ngomo, N. (2018).
\newblock {9th Challenge on Question Answering over Linked Data (QALD-9)}.
\newblock {\em language}, 7(1).

\bibitem[Nikas et~al., 2020]{nikas2020two}
Nikas, C., Fafalios, P., and Tzitzikas, Y. (2020).
\newblock Two-stage semantic answer type prediction for question answering
  using bert and class-specificity rewarding.
\newblock In {\em Proceedings of the SeMantic AnsweR Type prediction task
  (SMART 2020)}, pages 19--28.

\bibitem[Pan et~al., 2019]{pan2019entity}
Pan, J.~Z., Zhang, M., Singh, K., van Harmelen, F., Gu, J., and Zhang, Z.
  (2019).
\newblock Entity enabled relation linking.
\newblock In {\em International Semantic Web Conference}, pages 523--538.
  Springer.

\bibitem[Perevalov and Both, 2020]{perevalov2020augmentation}
Perevalov, A. and Both, A. (2020).
\newblock Augmentation-based answer type classification of the smart dataset.
\newblock In {\em Proceedings of the SeMantic AnsweR Type prediction task
  (SMART 2020)}, pages 1--9.

\bibitem[Rossiello et~al., 2021]{rossiello2021generative}
Rossiello, G., Mihindukulasooriya, N., Abdelaziz, I., Bornea, M., Gliozzo, A.,
  Naseem, T., and Kapanipathi, P. (2021).
\newblock Generative relation linking for question answering over knowledge
  bases.
\newblock In {\em International Semantic Web Conference}, pages 321--337.
  Springer.

\bibitem[Rybin et~al., 2021]{rubq2}
Rybin, I., Korablinov, V., Efimov, P., and Braslavski, P. (2021).
\newblock {RuBQ 2.0: An Innovated Russian Question Answering Dataset}.
\newblock {\em ESWC 2021}.

\bibitem[Saha et~al., 2018]{csqa2018}
Saha, A., Pahuja, V., Khapra, M.~M., Sankaranarayanan, K., and Chandar, S.
  (2018).
\newblock Complex sequential question answering: Towards learning to converse
  over linked question answer pairs with a knowledge graph.

\bibitem[Sakor et~al., 2019]{sakor2019old}
Sakor, A., Mulang, I.~O., Singh, K., Shekarpour, S., Vidal, M.~E., Lehmann, J.,
  and Auer, S. (2019).
\newblock Old is gold: linguistic driven approach for entity and relation
  linking of short text.
\newblock In {\em NAACL: HLT 2019}, pages 2336--2346.

\bibitem[Sakor et~al., 2020]{sakor2020falcon}
Sakor, A., Singh, K., Patel, A., and Vidal, M.-E. (2020).
\newblock Falcon 2.0: An entity and relation linking tool over wikidata.
\newblock In {\em Proceedings of the 29th ACM International Conference on
  Information \& Knowledge Management}, pages 3141--3148.

\bibitem[Setty and Balog, 2020]{setty2021semantic}
Setty, V. and Balog, K. (2020).
\newblock Semantic answer type prediction using bert: Iai at the iswc smart
  task 2020.
\newblock In {\em Proceedings of the SeMantic AnsweR Type prediction task
  (SMART 2020)}, pages 10--18.

\bibitem[Sorokin and Gurevych, 2018]{webspwd}
Sorokin, D. and Gurevych, I. (2018).
\newblock Modeling semantics with gated graph neural networks for knowledge
  base question answering.
\newblock In {\em Proceedings of the 27th International Conference on
  Computational Linguistics}, pages 3306--3317. Association for Computational
  Linguistics.

\bibitem[Steinmetz and Sattler, 2020]{steinmetz2020coala}
Steinmetz, N. and Sattler, K.-U. (2020).
\newblock Coala-a rule-based approach to answer type prediction.
\newblock In {\em Proceedings of the SeMantic AnsweR Type prediction task
  (SMART 2020)}, pages 29--40.

\bibitem[Trivedi et~al., 2017]{trivedi2017lc}
Trivedi, P., Maheshwari, G., Dubey, M., and Lehmann, J. (2017).
\newblock {LC-QuAD: A corpus for complex question answering over knowledge
  graphs}.
\newblock In {\em International Semantic Web Conference}, pages 210--218.
  Springer.

\bibitem[Usbeck et~al., 2019]{DBLP:journals/semweb/UsbeckRHCHNDU19}
Usbeck, R., R{\"{o}}der, M., Hoffmann, M., Conrads, F., Huthmann, J., Ngomo,
  A.~N., Demmler, C., and Unger, C. (2019).
\newblock Benchmarking question answering systems.
\newblock {\em Semantic Web}, 10(2):293--304.

\bibitem[Vallurupalli et~al., 2020]{vallurupalli2020fine}
Vallurupalli, S., Sleeman, J., Finin, T., et~al. (2020).
\newblock Fine and ultra-fine entity type embeddings for question answering.
\newblock In {\em Proceedings of the SeMantic AnsweR Type prediction task
  (SMART 2020)}, pages 57–--70.

\bibitem[Zhao et~al., 2015]{zhao2015self}
Zhao, H., Lu, Z., and Poupart, P. (2015).
\newblock Self-adaptive hierarchical sentence model.
\newblock In {\em Twenty-Fourth International Joint Conference on Artificial
  Intelligence}.

\bibitem[Zhou et~al., 2015]{zhou2015c}
Zhou, C., Sun, C., Liu, Z., and Lau, F. (2015).
\newblock {A C-LSTM neural network for text classification}.
\newblock {\em arXiv preprint arXiv:1511.08630}.

\end{thebibliography}

\documentclass{article}

\usepackage{arxiv}

\usepackage[utf8]{inputenc} 
\usepackage[T1]{fontenc}    
\usepackage{hyperref}       
\usepackage{url}            
\usepackage{booktabs}       
\usepackage{amsfonts}       
\usepackage{nicefrac}       
\usepackage{microtype}      
\usepackage{lipsum}
\usepackage{graphicx}
\usepackage{multirow}
\usepackage{listings}
\graphicspath{ {./images/} }

\usepackage{todonotes}

\title{Semantic Answer Type and \\Relation Prediction Task (SMART 2021)}

\author{
Nandana Mihindukulasooriya \\
IBM Research AI\\ 
  \texttt{nandana.m@ibm.com} \\
\And
  Mohnish Dubey\\
  University of Bonn \\and Fraunhofer IAIS\\
  \texttt{dubey@cs.uni-bonn.de} \\
\And
Alfio Gliozzo \\
IBM Research AI\\ 
  \texttt{gliozzo@us.ibm.com} \\
\And
Jens Lehmann \\
University of Bonn and Fraunhofer IAIS\\ 
\texttt{jens.lehmann@cs.uni-bonn.de} \\
\And
Axel-Cyrille Ngonga Ngomo \\
Universität Paderborn\\ 
\texttt{axel.ngonga@upb.de} \\
\And
Ricardo Usbeck \\
Conversational AI, Fraunhofer IAIS Dresden\\
\texttt{ricardo.usbeck@iais.fraunhofer.de} \\
\And
Gaetano Rossiello \\
IBM Research, USA\\
\texttt{gaetano.rossiello@ibm.com} \\ 
\And
Uttam Kumar \\
 University of Bonn, Germany\\
\texttt{uttam1216@gmail.com} \\
}

\begin{document}
\maketitle
\begin{abstract}
Each year the International Semantic Web Conference organises a set of Semantic Web Challenges to establish competitions that will advance state of the art solutions in some problem domains. The \textit{Semantic Answer Type and Relation Prediction Task} (SMART) task is one of the ISWC 2021 Semantic Web challenges. This is the second year of the challenge after a successful SMART 2020 at ISWC 2020. This year's version focuses on two sub-tasks that are very important to Knowledge Base Question Answering (KBQA): Answer Type Prediction and Relation Prediction. Question type and answer type prediction can play a key role in knowledge base question answering systems providing insights about the expected answer that are helpful to generate correct queries or rank the answer candidates. More concretely, given a question in natural language, the first task is, to predict the answer type using a target ontology (\textit{e.g.}, \textit{DBpedia} or \textit{Wikidata}. Similarly, the second task is to identify relations in the natural language query and link them to the relations in a target ontology. This paper discusses the task descriptions, benchmark datasets, and evaluation metrics. More information, please visit \url{https://smart-task.github.io/2021/}.

\end{abstract}

\keywords{Answer Type Prediction \and Relation Linking \and Question Answering \and ISWC Semantic Web Challenge}

\section{Background}

Question Answering (QA) is a popular task in Natural Language Processing and Information Retrieval. The goal of a QA engine is commonly to answer a natural language question based on information contained in (a potentially predefined set of) datasets. An instantiation of the QA task is reading comprehension, in which the expected answers can be either a segment of text or span from the corresponding reading passage of text. In contrast, Knowledge Graph Question Answering (KGQA), the answer is extracted from a structured knowledge graph using a query. Typically, these knowledge graphs are in RDF and the queries are formulated using the SPARQL query language. 

In KGQA, the expected answer can either be (1) a set of entities or literals from a given knowledge graph $G$ or (2) derived from an aggregation of facts found in $G$. Many available datasets for the KGQA task originate from the ISWC community (e.g., QALD~\cite{ngomo20189th}, LC-QuAD 1.0~\cite{trivedi2017lc}, LC-QuAD 2.0~\cite{dubey2019lc}, SimpleQuestions-DBpedia~\cite{azmy2018farewell}, SimpleQuestions-Wikidata~\cite{wikidata-benchmark}).

In KGQA, given a natural language query, the final goal is to formulate a logical query that can generate the answer to the questions. In SMART 2021, we will focus on two sub-tasks that are relevant to KBQA: (a) answer type prediction, and (b) relation prediction.  SMART 2021 is the second edition of the SMART task challenge (\url{https://smart-task.github.io/}). Its first incarnation happened at ISWC 2020 with 11 submitting systems in total, advancing the state of the art of answer type classification. The proceedings of the first edition is published in CEUR~\cite{DBLP:conf/semweb/2020smart} which consists of 8 submissions~\cite{perevalov2020augmentation, setty2021semantic, nikas2020two, steinmetz2020coala, ammar2020methodology, kertkeidkachorn2020hierarchical, vallurupalli2020fine, bill2020question}.

Classifying a question's answer type plays a key role in KGQA~\cite{harabagiu2000falcon,allam2012question}. The questions themselves can be generally classified based on Wh-terms (\textit{Who}, \textit{What}, \textit{When}, \textit{Where}, \textit{Which}, \textit{Whom}, \textit{Whose}, \textit{Why}). The answer type classification is the task of determining the type of the expected answer based on the query.  In literature, such answer type classifications is performed as a short-text classification task using a set of coarse-grained types, with either 6 types~\cite{zhao2015self,zhou2015c}
or fine-grained with 50 types~\cite{li2006learning} like in the TREC QA task\footnote{\url{https://trec.nist.gov/data/qamain.html}}. We propose that a more granular answer type classification is possible and usefule using popular Semantic Web ontologies such as DBepdia and Wikidata.

In addition to answer type prediction task which we also had last year, we will introduce the subtask of relation prediction \cite{rossiello2021generative,naseem2021semantics,earl,mihindukulasooriya2020leveraging,sakor2020falcon,lin2020kbpearl,sakor2019old,pan2019entity} as another step towards a functional KGQA system. This decision is also based on last year's systems performances and their implications. Relation prediction is a vital step towards formulating the correct query to extract the answer from the knowledge base. A question can contain one or more relations from an ontology connecting the entities and classes directly or indirectly mentioned in the question~\cite{DBLP:journals/semweb/UsbeckRHCHNDU19}. Participants will be able to choose to participate in one or both tasks using their systems. The two tasks will maintain two leaderboards.


\section{Task Description}
\subsection{Answer Type Prediction}
Given a natural language question, the task is to produce a ranked list of answer types of a given target ontology. Currently, the target ontology could be either \textit{DBpedia} or \textit{Wikidata}. Table~\ref{tab1:examples} illustrates some examples. The participating systems can be either supervised (training data is provided) or unsupervised.

\begin{table}[htb!]
\centering
\caption{Example questions and answer types}\label{tab1:examples}
\begin{tabular}{|p{7cm}|l|l|}
\hline
\multirow{2}{*}{\textbf{Question}}                   & \multicolumn{2}{c|}{\textbf{Answer Type}}                                      \\ \cline{2-3} 
                                                     & \multicolumn{1}{c|}{\textbf{DBpedia}} & \multicolumn{1}{c|}{\textbf{Wikidata}} \\ \hline
Give me all actors starring in movies directed by and starring William Shatner. & dbo:Actor & wd:Q33999 \\ \hline
Who is the heaviest player of the Chicago Bulls?     & dbo:BasketballPlayer                  & wd:Q3665646                            \\ \hline
How many employees does Google have?     & xsd:integer                  & xsd:integer                             \\ \hline
\end{tabular}
\end{table}

\subsection{Relation Prediction}
The second task is similar to the answer type prediction task. Here, the task is predict a set of relations from a given natural language question based on the underlying knowledge graph. The knowledge graph is either \textit{DBpedia} or \textit{Wikidata}. Table~\ref{tab2:examples} illustrates some examples. The participating systems can be either supervised (training data is provided) or unsupervised.
Note, an answer span detection is not needed nor do implicite relations from general ontologies such as \textit{rdf:type} or \textit{rdfs:label} need to be detected.


\begin{table}[htb!]
\centering
\caption{Example questions and relations}\label{tab2:examples}
\begin{tabular}{|p{6cm}|l|l|}
\hline
\multirow{2}{*}{\textbf{Question}}                   & \multicolumn{2}{c|}{\textbf{Relations}}                                      \\ \cline{2-3} 
                                                     & \multicolumn{1}{c|}{\textbf{DBpedia}} & \multicolumn{1}{c|}{\textbf{Wikidata}} \\ \hline
Give me all actors starring in movies directed by and starring William Shatner. & dbo:starring, dbo:director & wdt:P161, wdt:P57  \\ \hline
Which programming languages were influenced by Perl? & dbo:influencedBy               &              wdt:P737                  \\ \hline
How many employees does Google have?     &  dbo:numberOfEmployees                 &     wdt:P1128                         \\ \hline
\end{tabular}
\end{table}

\section{Datasets}

Rather than building a benchmark from scratch, several datasets for semantic answer type prediction are created from existing academic benchmarks 
for Knowledge Graph Question Answering. For creating the gold standards, we have used QALD-9~\cite{ngomo20189th}, LC-QuAD v1.0~\cite{trivedi2017lc}, and LC-QuAD v2.0~\cite{dubey2019lc} datasets. To overcome the problem of repeating last year's data, we have re-split the datasets and added new datasets. Tentatively, we are planing to add questions from RuBQ~\cite{rubq2}, SimpleQuestionsWikidata~\cite{wikidata-benchmark}, SimpleDBpediaQA~\cite{azmy2018farewell}, WebQSP-WD~\cite{webspwd}, and Complex Sequential Question Answering~\cite{csqa2018} datasets. Additionally, we have added a certain amount of honey pot questions to avoid cheating. 

\textbf{Answer Type Prediction: } 
Each of these datasets has a natural language query and a corresponding SPARQL query. We used the gold standard SPARQL query to generate results and analyze them to generate an initial answer type for each query and manually validated them. Table\ref{tab:ATData}

\begin{table}[h!]
\centering
\caption{Answer Type Prediction Datasets}\label{tab:ATData}
\begin{tabular}{|l|l|l|l|}
\hline
Dataset & Train  & Test   & Total  \\ \hline
SMART2021-AT-DBpedia      & 40,621 & 10,093 & 50,714 \\ \hline
SMART2021-AT-Wikidata       & 43,604 & 10,864 & 54,468 \\ \hline
\end{tabular}
\end{table}

\textbf{Relation Prediction:}
Similar to the answer type prediction task, for the relation prediction task, we extracted the corresponding relations from the given SPARQL queries. The datasets have natural language queries and their corresponding SPARQL queries. Each SPARQL query has a set of triple patterns, each containing a relation. For each natural language questions, those relations are considered as the gold standard relations.

\begin{table}[h!]
\centering
\caption{Relation Prediction Datasets}\label{tab:RLData}
\begin{tabular}{|l|l|l|l|}
\hline
Dataset & Train  & Test   & Total  \\ \hline
SMART2021-RL-DBpedia      & 34,204 & 8,552 & 42,756 \\ \hline
SMART2021-RL-Wikidata       & 24,112 & 6,029 & 30,141 \\ \hline
\end{tabular}
\end{table}


\section{Evaluation metrics and software}

The systems can participate in one or both tasks in one or both datasets. The are expected to provide a list of answer types or/and a set of relations from the target ontology for each natural language question. The answer types will be evaluated using the metrics: \textit{Mean Reciprocal Rank} (\textit{MRR}), \textit{Hits @ 1/5/10}, and \textit{Macro F1}. The systems will be ranked based on F1. The relations will be evaluated using \textit{Precision}, \textit{Recall}, \textit{Micro F1} and  \textit{Macro F1} and ranked according to the latter metric.

We provide evaluation scripts\footnote{\url{https://github.com/smart-task/smart-2021-dataset/tree/main/evaluation}} (in Python) for each task. 

\pagebreak

\bibliographystyle{apalike}  
\bibliography{ms}

\end{document}